OpenStudy (thomasbrewer98):

the rectangular part of the field below is 100 yards long and the diameter of each semicircle is 46 yards.Find the cost of reseeding the field at $0.60 per square yard how do you solve this so confuse

OpenStudy (anonymous):

1. Diameters of the 2 semicircles would form the edges of rectangular field. So use the diameter as the width to calculate rectangular area. 2. Then calculate the area of the 2 semi circles, radius 46/2. 3. Add both rectangular and semi circular areas 4. Multiply total area with $0.60 to find total cost.

OpenStudy (thomasbrewer98):

ok hold on

OpenStudy (thomasbrewer98):

what about the 100

OpenStudy (anonymous):

Yes 100 is the length of the rectangle. So, area of rectangular area = 100*46

OpenStudy (thomasbrewer98):

o ok

OpenStudy (thomasbrewer98):

so it is 40.100=? then u divide it by 2 then add it with the 0.60

OpenStudy (anonymous):

Total Area = Rectangular Area + 2x Semicircle Area = \[46\times100+2\times (\pi(46/2)^{2})/2\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (thomasbrewer98):

yes kinda

OpenStudy (thomasbrewer98):

so the answer will be 6261.90 ?

OpenStudy (anonymous):

Yes, thats your total area. Now you can find the total cost from the per sq. yard cost.

OpenStudy (thomasbrewer98):

now u add that with the 0.60

OpenStudy (anonymous):

Add? You sure about that?

OpenStudy (thomasbrewer98):

idk

OpenStudy (anonymous):

1 baseball bat costs $10. How much would it cost to by 10 baseball bats?

OpenStudy (anonymous):

See where I am going with this? You are almost there...just think how you will get total cost.

OpenStudy (thomasbrewer98):

mutiply

OpenStudy (anonymous):

there you go!

OpenStudy (thomasbrewer98):

ok so the answe is 3757.14